OpenStudy (anonymous):

|-2 2 4 | |-1 5-5| |4 -3 1 | evaluate the determinant I got -6

hartnn (hartnn):

i got -86

OpenStudy (anonymous):

ok, well both are options. I must have done my addition wrong

hartnn (hartnn):

show what u done ?

OpenStudy (anonymous):

I evaluated -10+(-40)+12-80-(-30)-(-2)

hartnn (hartnn):

first term -2 (5-15) = ... ?

OpenStudy (anonymous):

ok... I got -86 that time also haha

OpenStudy (anonymous):

-25?

hartnn (hartnn):

u got -86...thats correct...

OpenStudy (anonymous):

oh, I think I misunderstood your question though, on the first term (-2)(5)(1) I got -10

hartnn (hartnn):

but it should be +10

hartnn (hartnn):

+20

OpenStudy (anonymous):

-2, 5 and 1? how would that be positive?

hartnn (hartnn):

ohh..that way then -2(5*1) = -10 then -2(-(-5)(-3)) = +30 and -10+30 = 20 thats how i got +20

OpenStudy (anonymous):

oh, gotcha.

OpenStudy (anonymous):

do you work it differently then?

hartnn (hartnn):

lil bit...i solve it having 3 terms, instead of 6....basically same thing.

OpenStudy (anonymous):

oh ok, thats cool. I'd ask how, but I'm not looking to impress anyone, just making sure I got the right answer haha. I appreciate the assistance though.

hartnn (hartnn):

welcome ^_^